OpenStudy (anonymous):

PLease solve: (gºf)(x)=g(f(x)): 3(1/3x+2)-6

hero (hero):

I'm going to assume f(x) = 1/3x + 2 g(x) = 3x - 6 right?

OpenStudy (anonymous):

yes

hero (hero):

Okay, well distribute the 3 to finish simplifying

OpenStudy (anonymous):

then 3/3=0 3*x=3x 3*2=6 so then the equation would look something like this? 3x+6-6=3x?

hero (hero):

3/3 = 0? How?

hero (hero):

I thought 3/3 = 1

OpenStudy (jwheele1):

any number over its clone = 1. For example, 1/1=1; 2/2=1; 654/654=1

hero (hero):

The expression, if everything you wrote above is correct, should be: g(f(x)=x

OpenStudy (anonymous):

but then how would i graph this?

hero (hero):

Plug in values for x and solve for g(f(x)). My guess is you'll get something similar to the graph of y = x

OpenStudy (jwheele1):

This is a great resource by the way; http://www.khanacademy.org/math/algebra/quadtratics

hero (hero):

@jwheele1, I get the feeling that she's trying to figure out how to graph g(f(x))